# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 4.4.5 Harmonic representatives
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
@@
There exists $B_{i}$ $(i=1,2)$ such that the following holds on $X$ for any $t\geq 1$:
@@
Then, ${\tt V}(\tau,t)$ is a harmonic $1$-form satisfying $[{\tt V}(\tau,t)]=\iota_{t}^{{\tt v}}([\tau])$.
@@ 4.4.6 Estimate of the pairings
There exists $B_{i}$ $(i=1,2)$ such that
@@
$-2\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=-2\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}.$
Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we have
@@
Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}(\nu,t)=0$, we have

### v3

@@ 4.4.5 Harmonic representatives
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
@@
There exist $B_{i}$ $(i=1,2)$ such that the following holds on $X$ for any $t\geq 1$:
@@
Then, ${\tt V}(\tau,t)$ is a harmonic $1$-form satisfying ${\tt V}(\tau,t)]=\iota_{t}^{{\tt v}}([\tau])$.
@@ 4.4.6 Estimate of the pairings
There exist $B_{i}$ $(i=1,2)$ such that
@@
$2\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=-2\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}.$
Because $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we have
@@
Because $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}(\nu,t)=0$, we have